OpenStudy (anonymous):

Factor 5x^2 - 7x + 2. (5x + 2)(x - 1) (5x - 1)(x - 2) (5x - 2)(x - 1)

OpenStudy (anonymous):

(5x-2)(x-1)

OpenStudy (anonymous):

Factor 8y^2 + 6y + 1

OpenStudy (anonymous):

(4y+1)(2y+1)

OpenStudy (anonymous):

Find the length and width of a rectangle with an area of 2x^2 + x - 3. l = 2x + 3; w = x + 1 l = 2x + 3; w = x - 1 l = 3x + 2; w = x - 1